Run Meeting: August 20th, 2020

From clas12-run
Jump to: navigation, search

Back to Run Group F | Daily Run Group F Coordination Meetings

12:00PM RC Meeting (https://bluejeans.com/7576835804)

RC: M. Hattawy


2020

March
19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30,31
July
31
August
1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20 21 22 23 24, 25, 26, 27, 28, 29, 30, 31
September
1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20 21

Summary of Aug. 20th

Owl shift

  1. Shift summary by the shift expert: https://logbooks.jlab.org/entry/3831199
  2. Issues: communication issues with the sub-detectors of sector 4. Nathan Baltzell rebooted the icos and claimed that the might cam from beam radiation.
  3. Recorded data: 49M events on full D2 target.

Day shift

  1. Shift summary by the shift expert: https://logbooks.jlab.org/entry/3831473 ( normal production shift on full D2)
  2. Issues: DAQ crashed at 8 am. RTPC tripped once at 9:00 ( run 12614). Major vacuum leak in NL. No beam from 12:20 pm.
  3. Recorded data: 18M on full D2.

Swing shift

  1. Shift summary by the shift expert: https://logbooks.jlab.org/entry/3831666
  2. Issues: PD notified that the situation got worse when they tried to send beam to halls. there will be no beam till an expert check it tomorrow at 10 am. The owl shift tonight is cancelled.
  3. Recorded data: 0

Whole day summary

  1. Total recorded data today: 67M events.
  2. Part-II total accumulated data on D2: 754M (25%).
  3. Concerns: We still need to solve the puzzle of the increased scalers and DC occupancy with the new trigger file when we get beam back.


Tomorrow daily RC meeting will be at usual time 12:00 PM using the link https://bluejeans.com/7576835804/

Recorded data numbers

Number of events recorded